TB MC Qu. 25-104 Rosies Company has three products... Rosies Company has three products, P1, P2, and P3. The maximum Rosie
Answer #1
P1 P2 P3
Contribution margin per unit 5 15 25
X Units per hour 12 6 3
Contribution margin per hour 60 90 75
Company should produce P2 first, then P3 and at last P1
Units of P2 = 24000
Hours for P2 = 24000/6= 4000
Units of P3 = 12000
Hours for P3 = 12000/3= 4000
Remaining hours for P1 1000 =9000-4000-4000
Units of P1 = 1000*12 = 12000
12,000 P1, 24,000 P2, 12,000 P3
Option A is correct